Suppose a 350-g kookaburra (a large kingfisher bird) picks up a 75-g snake and raises it 2.5 m from the ground to a branch.

(a) How much work did the bird do on the snake?

(b) How much work did it do to raise its own center of mass to the branch?


1
Expert's answer
2021-03-10T17:21:43-0500

(a)

"W_s=F_sh=m_sgh=0.075\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot2.5\ m=1.84\ J."

(b)

"W_b=F_bh=m_bgh=0.35\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot2.5\ m=8.6\ J."